# Unveiling Interpretability in Self-Supervised Speech Representations for Parkinson’s Diagnosis

### v1

@@ Abstract
However, despite the improved performance, SSL-based speech representations present notable limitations in interpretability. Due to their complex, often opaque structure, understanding how these models capture specific speech characteristics — such as phonatory, articulatory, or prosodic features — remains particularly challenging. In clinical settings, however, this black-box nature of SSL models raises important concerns, as diagnostic support systems must be interpretable and explainable to promote trust, transparency, and compliance with ethical and legal standards [rasheed2022explainable]. While recent studies have started to explore interpretability in SSL models broadly [pasad2021layerwise, pasad2023comparative, singla2022whatdo], research specifically tailored to pathological speech is still quite limited [baumann2024cleftlip, escobar2024foundation, gallo2024levels]. In this line, although there is an ongoing debate regarding the reliability of attention mechanisms for interpretability [jain2019not, wiegreffe2019notnot, bibal2022debate], attention-based model interpretations have recently shown promising achievements in studies on cleft lip and palate pathological speech [baumann2024cleftlip, baumann24cleftattn]. Indeed, these methods not only provided interpretable insights, but also present a more resource-efficient alternative to their gradient-based counterparts, making them especially appealing for practical use in clinical settings.
@@ III-B Interpretability through Cross-Attention Mechanisms
To achieve this, our proposed model uses both SSL-based speech embedding sequences and informed speech feature sets as inputs. Specifically, each utterance sample is represented by an SSL embedding sequence, , which captures temporal speech patterns in the data, along with an informed feature set, , which contains more interpretable but static speech characteristics. Our goal is therefore to obtain two attention score matrices that reflect the relationships between both speech representations. The first attention score matrix, , would address interpretability at the embedding level, where for each dimension of the SSL-based embedding, we quantify the contribution of each of the knowledge-based, informed features. Similarly, we aim to obtain a second attention score matrix, , which would provide a complementary temporal-level perspective, providing the contribution of each of the informed features at each time step of the SSL sequence.
@@
where denote learnable linear projections, while corresponds to the identity matrix , representing a non-transformative operation that leaves the informed features unchanged. This choice, along with the decision to preserve the dimensionality of the SSL-based representations intact, were made intentionally to support our goal of interpretability within the proposed framework. Therefore, in general terms, the informed speech features act directly as the key values, providing a set of interpretable dimensions against which SSL-based embeddings are aligned. However, our dual-branch architecture was designed to investigate model interpretability from two distinct, yet complementary, perspectives. Here is where the static nature of our informed speech features plays a crucial role, offering flexibility to expand them alongside either the temporal or embedding feature dimensions. Depending on the interpretability perspective addressed, different aspects, including slight modifications of the original attention mechanism — specifically matrix translations to align shapes for the derived dot products — were considered:
@@ IV-A Self-Supervised Speech Features
Finally, existing literature on layer-wise analysis of SSL-based foundational speech models [pasad2021layerwise] suggests that these architectures exhibit an autoencoder-like behavior, with deeper layers becoming progressively more abstract, while the final layers start to resemble the input speech features, almost as if reconstructing them. These same studies also found that shallow, intermediate layers tend to capture more acoustic, phonetic, and linguistic information, which aligns with the findings of our preliminary experiments and supports the choice of the 7th layer as an effective representation for interpretability in the context of PD diagnosis. Similar conclusions were reported by Favaro et al. [favaro2023interpretable].
IV-B Interpretable Speech Features

### v2

@@ Abstract
However, despite the improved performance, SSL-based speech representations present notable limitations in interpretability. Due to their complex, often opaque structure, understanding how these models capture specific speech characteristics — such as phonatory, articulatory, or prosodic features — remains particularly challenging. In clinical settings, this black-box nature of SSL models raises important concerns, as diagnostic support systems must be interpretable and explainable to promote trust, transparency, and compliance with ethical and legal standards [rasheed2022explainable]. While recent studies have started to explore interpretability in SSL models broadly [pasad2021layerwise, pasad2023comparative, singla2022whatdo], research specifically tailored to pathological speech is still quite limited [baumann2024cleftlip, escobar2024foundation, gallo2024levels]. In this line, although there is an ongoing debate regarding the reliability of attention mechanisms for interpretability [jain2019not, wiegreffe2019notnot, bibal2022debate], attention-based model interpretations have recently shown promising achievements in studies on cleft lip and palate pathological speech [ariasvergara2023measuring, baumann2024cleftlip, baumann24cleftattn]. Indeed, these methods not only provided interpretable insights, but also present a more resource-efficient alternative to their gradient-based counterparts, making them especially appealing for practical use in clinical settings.
@@ III-B Interpretability through Cross-Attention Mechanisms
To achieve this, our proposed model uses both SSL-based speech embedding sequences and informed speech feature sets as inputs. Specifically, each utterance sample is represented by an SSL embedding sequence, , which captures temporal speech patterns in the data, along with an informed feature set, , which contains more interpretable but static speech characteristics. Here, denotes the number of time steps, is the embedding dimension, and represents the number of clinically informed speech features.
Our goal is therefore to obtain two attention score matrices that reflect the relationships between both speech representations. The first attention matrix, , would address interpretability at the embedding level, where for each dimension of the SSL-based embedding, we quantify the contribution of each of the knowledge-based, informed features. Similarly, we aim to obtain a second attention matrix, , which would offer a complementary temporal-level perspective, providing the contribution of each of the informed features at each time step of the SSL sequence.
@@
where denote learnable linear projections, while corresponds to the identity matrix , representing a non-transformative operation that leaves the informed features unchanged. This choice, along with the decision to preserve the dimensionality of the SSL-based representations intact, was made intentionally to support our goal of interpretability within the proposed framework. Therefore, in general terms, the clinically informed speech features act directly as the key values, providing a set of interpretable dimensions against which SSL-based embeddings are aligned. However, our dual-branch architecture was designed to investigate model interpretability from two distinct, yet complementary, perspectives. Here is where the static nature of our informed speech features plays a crucial role, offering flexibility to expand them alongside either the temporal or embedding feature dimensions. Depending on the interpretability perspective addressed, different aspects, including slight modifications of the original attention mechanism — specifically matrix translations to align shapes for the derived dot products — were considered:
@@ IV-A Self-Supervised Speech Features
Finally, existing literature on layer-wise analysis of SSL-based foundational speech models [pasad2021layerwise] suggests that these architectures exhibit an autoencoder-like behavior, with deeper layers becoming progressively more abstract, while the final layers start to resemble the input speech features, almost as if reconstructing them. These same studies also found that shallow, intermediate layers tend to capture more acoustic, phonetic, and linguistic information. Consistent with these findings, recent works in the context of pathological speech analysis [favaro2023interpretable, purohit2025layeradapt] have demonstrated that intermediate layer representations can achieve performance comparable to fine-tuning the entire encoder architecture, thus providing a more parameter-cost efficient alternative. Our preliminary layer-wise analysis also revealed that using the 7th layer consistently outperformed using other layers, by a larger margin when compared to extreme layers (e.g., the first and last), and by a smaller margin to adjacent layers. In conclusion, all these findings further support our approach as an effective representation for interpretability in the context of PD diagnosis.
IV-B Interpretable Clinically Informed Speech Features